OpenStudy (anonymous):

x^2 + 10x + 21/x^3 + x^2 - 42x Simplify

OpenStudy (anonymous):

you need to factor each (x+7)(x+3) is the numerator x(x^2+x+42) can be factored more as the denominator and then you can cancel a term

OpenStudy (anonymous):

Think you could simplify the bottom more I am still confused about the bottom really the only thing I was struggling with...

OpenStudy (anonymous):

x(x+7)(x-6) it is -42 not +42 that I wrote

OpenStudy (anonymous):

Oh so you factor out the 42 again as if it was like the numerator but leave the x on front???

OpenStudy (anonymous):

x^3+x^2-42x=x(x^2+x-42)=x(x+7)(x-6) then the x-7 s cancel and you are left with \[\frac{ x+3 }{ x(x-6) }\]

OpenStudy (anonymous):

Ok thank you so much!